3a. How many grams of sucrose must be added to 430g of water to lower the vapor pressure 1.4mmHg at 25C? The vapor pressure of water at 25C is 233.8 mmHg and molar mass is 342.3 g/mol3. Vapor pressure lowering a) How many grams of sucrose must be added to 430 R of water to lower the vapor pressure 1.4 mm Hg
Answer #1

3a)

No. of moles of sucrose = n1

No. of moles of water = n2 = (430 g)/(18.02 g/mole) = 23.86 mole

Vapor pressure lowering = Δp= 1.4 mmHg

Vapor pressure of pure water = p0 = 23.8 mmHg

Calculate the no. of moles of sucrose using following formula

(Δp/p0) = (n1)/( n1+ n2)

(1.4/23.8) = (n1)/( n1+ 23.86 mole)

(0.0588)(n1 + 23.86 mole) = n1

0.0588 n1 + 1.403 = n1

n1 - 0.0588 n1 = 1.403

0.941n1 = 1.403

n1 = 1.403/0.941 = 1.491 mole

No. of moles of sucrose = 1.491 mole

Grams of sucrose = (No. of moles of sucrose)(Molar mass of sucrose)

= (1.491 mole)(342.3 g/mole)

= 510.36 g

Grams of sucrose = 510.4 g